# CTP-SCU/2025023

### v1

@@ 1 Introduction
If Susskind and Uglum’s conjecture is correct, there must exist a correspondence between the string worldsheet and the RT surface. Such a correspondence could plausibly serve as a guiding principle for developing the open string perspective of the conjecture in the framework of bosonic string theory, ultimately enabling the computation of entanglement entropy using the string worldsheet. With this framework in place, the black hole information paradox can be revisited from a new perspective, potentially revealing the microscopic quantum state of Hawking radiation. In previous works, qualitative evidence for a relationship between the string worldsheet and the RT surface has been found in both open and closed string field theories [13, 14]. Both entanglement entropy and string field theory are fundamentally grounded in hyperbolic geometry. Furthermore, key dynamical features—such as phase transitions in the entanglement wedge cross section (EWCS) [15] or reflected surface [16], and the Batalin–Vilkovisky (BV) master equations in open or closed string field theory [17, 18]—appear to be intrinsically linked. These observations suggest that the correspondence is not a mere coincidence due to shared geometric structures but reflects a deeper, underlying connection. Nevertheless, a rigorous quantitative proof of this correspondence remains elusive. Moreover, as emphasized in [11], previous attempts to compute entanglement entropy using string theory [19, 20] suffer from a conceptual problem. In those approaches, the “position” of a string is defined solely by the location of its center of mass. However, due to string vibrations, different parts of the string can extend outside the assigned region, making the center-of-mass prescription inconsistent with the actual spatial support of the string.
@@
We define the partial derivatives of $\Omega$ with respect to $X$ and $X^{\prime}$ as $\Omega_{\mu}\coloneqq\partial\Omega/\partial X^{\mu}$ and $\Omega_{\mu^{\prime}}\coloneqq\partial\Omega/\partial X^{\prime\mu}$, respectively. The covariant derivative of $\Omega_{\mu}$ with respect to $X$ is denoted as $\Omega_{\mu\nu}\coloneqq\nabla_{\nu}\Omega_{\mu}$, while the mixed derivative with respect to $X^{\prime}$ is $\Omega_{\mu\nu^{\prime}}\coloneqq\partial_{\nu^{\prime}}\Omega_{\mu}$. An important concept in this context is the coincidence limit, in which $X^{\prime}\rightarrow X$, denoted by $\left[\cdots\right]$. Under this limit, the spacetime metric derived from the geodesic length is given by
@@
Here, a key insight emerges. In string theory, gravity is known to arise from CFT. When a string propagates in a curved background, the requirement that the worldsheet theory remain conformally invariant at the quantum level imposes constraints on the background geometry—namely, the string-theoretic version of the Einstein equations [23]. This observation raises a compelling question: can we apply the same logic to entanglement entropy using the expansion of $\Omega_{\mu\nu}$? More precisely, can we impose conformal symmetry on this expansion and derive Einstein equations? This question is well-motivated, especially since $\Omega_{\mu\nu}$, which encodes geometric data such as the geodesic separation between two points in spacetime, is entirely constructed from CFT correlators. If $\Omega_{\mu\nu}$ arises from CFT, then requiring quantum-level conformal invariance may indeed constrain the emergent geometry—potentially leading to gravitational dynamics from entanglement itself.
@@
$S_{E}=\frac{1}{\ell^{2}}\int d^{2}\sigma\Omega_{\mu\nu}\partial_{\alpha}\Omega^{\mu}\partial^{\alpha}\Omega^{\nu},\qquad\alpha=\left(\tau,\sigma\right).$ (6)
where $\ell$ denotes a characteristic length scale. We use the prime notation here because, by expanding the action in the near-coincidence limit $X\rightarrow X^{\prime}$ and employing RNC $\mathbb{X}$, we obtain
@@ 2 Motivation
[…]
@@
The partial derivatives of $\Omega$ with respect to $X$ and $X^{\prime}$ are denoted as $\Omega_{\mu}\coloneqq\partial\Omega/\partial X^{\mu}$ and $\Omega_{\mu^{\prime}}\coloneqq\partial\Omega/\partial X^{\prime\mu}$, respectively. Furthermore, the covariant derivative of $\Omega_{\mu}$ with respect to $X$ is given by $\Omega_{\mu\nu}\coloneqq\nabla_{\nu}\Omega_{\mu}$, while its derivative with respect to $X^{\prime}$ is $\Omega_{\mu\nu^{\prime}}\coloneqq\partial_{\nu^{\prime}}\Omega_{\mu}$. With these definitions, we introduce an important limit: the coincidence limit, which refers to the limit $X^{\prime}\rightarrow X$, denoted as $\left[\cdots\right]$. Under this limit, the spacetime metric can be expressed as
@@
However, taking this limit eliminates valuable information. Remarkably, it is possible to extract spacetime information without imposing the coincidence limit. Instead, we can expand this quantity in the near-coincidence regime $X^{\prime}\rightarrow X$:
$\Omega_{\mu\nu}\left(X,X^{\prime}\right)=g_{\mu\nu}\left(X\right)-\frac{1}{3}R_{\mu\lambda\nu\kappa}\left(X\right)\Omega^{\lambda}\left(X,X^{\prime}\right)\Omega^{\kappa}\left(X,X^{\prime}\right)+\ldots.$ (29)
This expansion closely resembles a procedure in string theory. As previously reviewed, requiring the worldsheet theory to be invariant under conformal transformations leads to the Einstein equations. Here, we pose a similar question: Can we impose conformal symmetry to obtain the Einstein equations? This is a reasonable question, especially since the quantity $\Omega_{\mu\nu}$ is entirely derived from CFT data. If $\Omega_{\mu\nu}$, encoding geometric information such as the geodesic separation between two spacetime points, emerges from CFT correlators, then requiring conformal invariance at the quantum level may constrain the background geometry. The challenge lies in how to impose conformal symmetry on a known quantity such as the geodesic length. Fortunately, such a possibility exists. The geodesic length in AdS3 can be related to the entanglement entropy of a CFT2 via the RT formula, which is inherently conformally invariant:
@@
At this step, the AdS3/CFT2 correspondence enters the story. It is worth noting that this requirement also ensures that the theory remains conformally invariant after quantization. The key point is that the geodesic length is a classical quantity, while entanglement entropy arises from the quantum theory. The RT formula serves as a bridge between these two regimes, relating a classical geometric object in the bulk to a quantum information-theoretic quantity in the boundary theory. Therefore, maintaining conformal invariance across both sides is essential for the consistency of the correspondence. Based on this argument, if we assume that there exists an action for entanglement entropy expressed in terms of $\Omega_{\mu\nu}\left(X\right)$, we can substitute the entanglement entropy expression (30) into the expansion (29), which leads to the following form of the $\beta$-function for the spacetime metric $g$:
@@ 3 The worldsheet action for entanglement entropy
Note $\left(t_{i},x_{i}\right)$ are ordinary coordinate of CFT2. Now, we can regroup the coordinates as follows [36] :
@@
$S_{\mathrm{vN}}\left(A:B\right)=\frac{c}{6}\cosh^{-1}\left(1+\frac{\triangle Z^{2}+\triangle X^{2}-\triangle T^{2}}{2ZZ^{\prime}}\right),$ (36)
@@
The situation changes if the entangling regions are varied in time. In this case, both the entanglement entropy and its dual geodesic length evolve accordingly. As a result, the geodesic effectively sweeps through AdS3, as illustrated in figure (6). The evolving trajectory generates a two-dimensional surface that can be described by an embedding function $X^{\mu}\left(\tau,\sigma\right)$. In other words, By promoting the two bulk endpoints $X^{\mu}$ and $X^{\mu\prime}$ to fields $X^{\mu}\left(\tau,\sigma\right)$, $X^{\mu\prime}\left(\tau,\sigma\right)$ we obtain a two-parameter family of geodesics. The union (envelope) of this family defines an emergent two-dimensional surface in the bulk. Motivated by the analogy with string theory, one can then construct an action for this surface—analogous to the Polyakov action—that captures the dynamics of entanglement entropy as the entangling regions change.
@@
where $\partial_{\alpha}=\left(\partial_{\tau},\partial_{\sigma}\right)$ acts only on the field point $X\left(\tau,\sigma\right)$. Both $X$ and $X^{\prime}$ are boundary-determined bulk points fixed by the entangling regions in the CFT. Since the two endpoints $X$ and $X^{\prime}$ of any geodesic are interchangeable, there exists a corresponding set of derivatives $\partial_{\alpha^{\prime}}=\left(\partial_{\tau^{\prime}},\partial_{\sigma^{\prime}}\right)$ acting on $X^{\prime}\left(\tau^{\prime},\sigma^{\prime}\right)$. As a result, we obtain a non-local action constructed from the entanglement entropy, which can describe the physics at either of the spacetime coordinates $X$ or $X^{\prime}$. To verify the consistency of this action, one must check whether it reproduces the entanglement entropy or geodesic length. Before this verification, however, let us examine its physical meaning. Here, instead of treating $X^{\mu}\left(\tau,\sigma\right)$ as the embedding of a true string, the embedding here is constructed out of geodesics in AdS that are dual to entangling intervals in the CFT2. The Synge’s world function $\Omega\left(X,X^{\prime}\right)$ (half the squared geodesic distance) replaces the role of $X^{\mu}$ as the basic dynamical variable. Consequently, the action is built directly from entanglement entropy data (through $\Omega$) rather than fundamental fields. The coordinates $\left(\tau,\sigma\right)$ parametrize the varying entangling regions. As these regions evolve, their dual geodesics sweep out a two-dimensional surface in AdS. This surface is not a fundamental string worldsheet but an effective entanglement worldsheet: a bookkeeping device that tracks the dynamics of entanglement entropy. The action (40) therefore encodes the dynamics of geodesic lengths (equivalently entanglement entropies) as the entangling regions change. The quadratic form in derivatives of $\Omega^{\mu}$ resembles the Polyakov action, ensuring that the dynamics is governed by variations of geodesic distance.
@@
where $g_{\mathrm{ab}}\mathbb{X}^{\mathrm{a}}\mathbb{X}^{\mathrm{b}}$ denotes the squared geodesic distance between the base point $X^{\prime}$ and the field point $X$. A key feature of RNC is that they shift the base point $X^{\prime}$ to the origin, so that $\mathbb{X}^{\mathrm{a}}=0$. On the other hand, since the Synge’s world function $\Omega\left(X,X^{\prime}\right)$ depends on both the base point $X^{\prime}$ and the field point $X$, and since these points can be parameterized by the boundary CFT2 coordinate $\left(\tau,\sigma\right)$, we can write $\Omega\left(X,X^{\prime}\right)=\Omega\left(X\left(\tau,\sigma\right),X^{\prime}\left(\tau^{\prime},\sigma^{\prime}\right)\right)$. Consequently, the RNC coordinates $\mathbb{X}^{\mathrm{a}}$ can also be expressed in terms of the CFT2 coordinate $\left(\tau,\sigma\right)$ as
@@
$+\frac{1}{\ell^{2}}\int_{\partial\Sigma}d\tau\left[2n_{\alpha^{\prime}}\Omega_{\mu^{\prime}\nu^{\prime}}\partial_{\rho^{\prime}}\Omega^{\mu^{\prime}}\partial^{\alpha^{\prime}}\Omega^{\nu^{\prime}}\delta X^{\rho^{\prime}}\right]_{\sigma=\mathrm{bdy}},$
@@
which is precisely the geodesic equation for the endpoints along the D-brane. This result can be interpreted as shown in figure (7). In our setup, the spacetime point $X$, representing a geodesic endpoint, is parameterized by two-dimensional coordinates $\left(\tau,\sigma\right)$. Consequently, the geodesic sweeps out a two-dimensional surface whose dynamics are governed by the non-local EOM and boundary conditions (50). Taking the near-coincidence limit $X^{\prime}\rightarrow X$ reduces these to the local EOM and boundary conditions (52), in full agreement with string theory. Under the appropriate boundary conditions, the boundary $\partial\Sigma$ obeys the geodesic EOM (53). Thus, the proposed action (40) indeed contains the expected geodesic dynamics.
Figure 7: The left panel shows how a family of geodesics between $X$ and $X^{\prime}$ (red lines) sweeps out an effective two-dimensional surface. The length of each geodesic corresponds to the bulk dual of a CFT entanglement entropy. In the near-coincidence limit $X^{\prime}\rightarrow X$, this surface acquires the structure of a string worldsheet, with the Polyakov action emerging naturally. The right panel illustrates this worldsheet. Depending on the gauge choice, the open string (blue lines) can fluctuate along horizontal or vertical directions, while its endpoints obey the geodesic equation for a point-like particle.
@@
$S_{E}=\frac{1}{\ell^{2}}\int d^{2}\sigma^{\prime}\sqrt{\gamma}\left(\gamma^{\alpha\beta}\Omega_{\mu^{\prime}\nu^{\prime}}\left(X,X^{\prime}\right)+i\epsilon^{\alpha\beta}\mathcal{B}_{\mu^{\prime}\nu^{\prime}}\left(X,X^{\prime}\right)\right)\partial_{\alpha^{\prime}}\Omega^{\mu^{\prime}}\left(X,X^{\prime}\right)\partial_{\beta^{\prime}}\Omega^{\nu^{\prime}}\left(X,X^{\prime}\right).$ (56)
@@
$\mathcal{B}_{\mu^{\prime}\nu^{\prime}}\left(X\right)$ $=$ $B_{\mu^{\prime}\nu^{\prime}}\left(X\right)+\nabla_{\lambda^{\prime}}B_{\mu^{\prime}\nu^{\prime}}\Omega^{\lambda^{\prime}}\left(X,X^{\prime}\right)+\frac{1}{2}\left(\nabla_{\alpha^{\prime}}\nabla_{\beta^{\prime}}B_{\mu^{\prime}\nu^{\prime}}\left(X\right)-\frac{1}{3}R_{\;\alpha^{\prime}\mu^{\prime}\beta^{\prime}}^{\lambda^{\prime}}B_{\lambda^{\prime}\nu^{\prime}}\right$. (57)

### v2

@@ 1 Introduction
If Susskind and Uglum’s conjecture is correct, there must exist a correspondence between the string worldsheet and the RT surface. Such a correspondence could plausibly serve as a guiding principle for developing the open string perspective of the conjecture in the framework of bosonic string theory, ultimately enabling the computation of entanglement entropy using the string worldsheet. With this framework in place, the black hole information paradox can be revisited from a new perspective, potentially revealing the microscopic quantum state of Hawking radiation. In previous works, qualitative evidence for a relationship between the string worldsheet and the RT surface has been found in both open and closed string field theories [13, 14]. Both entanglement entropy and string field theory are fundamentally grounded in hyperbolic geometry. Furthermore, key dynamical features—such as phase transitions in the entanglement wedge cross section (EWCS) [15] or reflected surface [16], and the Batalin–Vilkovisky (BV) master equations in open or closed string field theory [17, 18]—appear to be intrinsically linked. These observations suggest that the correspondence is not a mere coincidence due to shared geometric structures but reflects a deeper, underlying connection. Nevertheless, a rigorous quantitative proof of this correspondence remains elusive. A similar observation was recently made in [19]: as expected, the RT surface for multiboundary black holes—which in our framework corresponds to the string vertices discussed earlier—can be derived directly from the dual CFT2.
Moreover, there are several related approaches to computing entanglement entropy using string theory. The first method involves evaluating the entropy in string theory via orbifolds [20, 21, 22, 23, 24], where the orbifold fixed point corresponds to the tip of a cone with opening angle $\beta=2\pi/N$. However, the methods in [22, 23, 24] face a notable difficulty: they fail to reproduce the entropy in the $N=1$ limit, which is required for consistency with the Susskind–Uglum conjecture. On the other hand, as emphasized in [11], another approach [25, 26] suffers from a conceptual issue. In this approach, the “position” of a string is defined solely by the location of its center of mass. However, due to string vibrations, different parts of the string can extend outside the assigned region, making the center-of-mass prescription inconsistent with the actual spatial support of the string.
@@
We define the partial derivatives of $\Omega$ with respect to $X$ and $X^{\prime}$ as $\Omega_{\mu}\coloneqq\partial\Omega/\partial X^{\mu}$ and $\Omega_{\mu^{\prime}}\coloneqq\partial\Omega/\partial X^{\mu^{\prime}}$, respectively. The covariant derivative of $\Omega_{\mu}$ with respect to $X$ is denoted as $\Omega_{\mu\nu}\coloneqq\nabla_{\nu}\Omega_{\mu}$, while the mixed derivative with respect to $X^{\prime}$ is $\Omega_{\mu\nu^{\prime}}\coloneqq\partial_{\nu^{\prime}}\Omega_{\mu}$. An important concept in this context is the coincidence limit, in which $X^{\prime}\rightarrow X$, denoted by $\left[\cdots\right]$. Under this limit, the spacetime metric derived from the geodesic length is given by
@@
Here, a key insight emerges. In string theory, gravity is known to arise from CFT. When a string propagates in a curved background, the requirement that the worldsheet theory remain conformally invariant at the quantum level imposes constraints on the background geometry—namely, the string-theoretic version of the Einstein equations [29]. This observation raises a compelling question: can we apply the same logic to entanglement entropy using the expansion of $\Omega_{\mu^{\prime}\nu^{\prime}}$? More precisely, can we impose conformal symmetry on this expansion and derive Einstein equations? This question is well-motivated, especially since $\Omega_{\mu^{\prime}\nu^{\prime}}$, which encodes geometric data such as the geodesic separation between two points in spacetime, is entirely constructed from CFT correlators. If $\Omega_{\mu^{\prime}\nu^{\prime}}$ arises from CFT, then requiring quantum-level conformal invariance may indeed constrain the emergent geometry—potentially leading to gravitational dynamics from entanglement itself.
@@
$S_{E}=\frac{1}{\ell^{2}}\int d^{2}\sigma^{\prime}\Omega_{\mu^{\prime}\nu^{\prime}}\partial_{\alpha^{\prime}}\Omega^{\mu^{\prime}}\partial^{\alpha^{\prime}}\Omega^{\nu^{\prime}},\qquad\alpha^{\prime}=\left(\tau^{\prime},\sigma^{\prime}\right).$ (6)
where $\ell$ denotes a charactestic length scale. We use the prime notation here because, by expanding the action in the near-coincidence limit $X\rightarrow X^{\prime}$ [^1] Since $X$ and $X^{\prime}$ are interchangeable, one may also drop the prime in the action (6) and take the limit $X^{\prime}\rightarrow X$. and employing RNC $\mathbb{X}$, we obtain
@@ 2 Motivation
On the other hand, let us examine how a similar structure arises in the context of entanglement entropy. To proceed, we introduce Synge’s world function $\Omega$ [27]. The good review of Synge’s world function can be found in ref. [28]. Consider a geodesic $z^{\mu}\left(\tau\right)$ that connects two points in spacetime: the base point $X^{\prime}$ and the field point $X$. The affine parameter $\tau$ runs from $\tau_{0}$ to $\tau_{1}$, such that:
@@
The partial derivatives of $\Omega$ with respect to $X$ and $X^{\prime}$ are denoted as $\Omega_{\mu}\coloneqq\partial\Omega/\partial X^{\mu}$ and $\Omega_{\mu^{\prime}}\coloneqq\partial\Omega/\partial X^{\mu^{\prime}}$, respectively. Furthermore, the covariant derivative of $\Omega_{\mu}$ with respect to $X$ is given by $\Omega_{\mu\nu}\coloneqq\nabla_{\nu}\Omega_{\mu}$, while its derivative with respect to $X^{\prime}$ is $\Omega_{\mu\nu^{\prime}}\coloneqq\partial_{\nu^{\prime}}\Omega_{\mu}$. With these definitions, we introduce an important limit: the coincidence limit, which refers to the limit $X^{\prime}\rightarrow X$, denoted as $\left[\cdots\right]$. Under this limit, the spacetime metric can be expressed as
@@
However, taking this limit eliminates valuable information. Remarkably, it is possible to extract spacetime information without imposing the coincidence limit. Instead, we can expand this quantity in the near-coincidence regime $X\rightarrow X^{\prime}$:
This expansion closely resembles a procedure in string theory. As previously reviewed, requiring the worldsheet theory to be invariant under conformal transformations leads to the Einstein equations. Here, we pose a similar question: Can we impose conformal symmetry to obtain the Einstein equations? This is a reasonable question, especially since the quantity $\Omega_{\mu^{\prime}\nu^{\prime}}$ is entirely derived from CFT data. If $\Omega_{\mu^{\prime}\nu^{\prime}}$, encoding geometric information such as the geodesic separation between two spacetime points, emerges from CFT correlators, then requiring conformal invariance at the quantum level may constrain the background geometry. The challenge lies in how to impose conformal symmetry on a known quantity such as the geodesic length. Fortunately, such a possibility exists. The geodesic length in AdS3 can be related to the entanglement entropy of a CFT2 via the RT formula, which is inherently conformally invariant:
@@
At this step, the AdS3/CFT2 correspondence enters the story. It is worth noting that this requirement also ensures that the theory remains conformally invariant after quantization. The key point is that the geodesic length is a classical quantity, while entanglement entropy arises from the quantum theory. The RT formula serves as a bridge between these two regimes, relating a classical geometric object in the bulk to a quantum information-theoretic quantity in the boundary theory. Therefore, maintaining conformal invariance across both sides is essential for the consistency of the correspondence. Based on this argument, if we assume that there exists an action for entanglement entropy expressed in terms of $\Omega_{\mu^{\prime}\nu^{\prime}}\left(X^{\prime}\right)$, we can substitute the entanglement entropy expression (30) into the expansion (29), which leads to the following form of the $\beta$-function for the spacetime metric $g$:
@@ 3 The worldsheet action for entanglement entropy
Note $\left(t_{i},x_{i}\right)$ are ordinary coordinate of CFT2. We may regroup these coordinates into effective bulk coordinates $\left(T,X,Z\right)$. The specific regrouping transformation is given in [42, 43, 44], while the general transformation was derived in [14] and reads
@@
$S_{\mathrm{vN}}\left(A:B\right)=\frac{c}{6}\cosh^{-1}\left(1+\frac{\triangle Z^{2}+\triangle X^{2}-\triangle T^{2}}{2ZZ^{\prime}}\right).$ (36)
@@
The situation changes if the entangling regions are varied in time. In this case, both the entanglement entropy and its dual geodesic length evolve accordingly. As a result, the geodesic effectively sweeps through AdS3, as illustrated in figure (6). The evolving trajectory generates a two-dimensional surface that can be described by an embedding function $X^{\mu}\left(\tau,\sigma\right)$. In other words, By promoting the two bulk endpoints $X^{\mu}$ and $X^{\mu\prime}$ to fields $X^{\mu}\left(\tau,\sigma\right)$, $X^{\mu\prime}\left(\tau^{\prime},\sigma^{\prime}\right)$ we obtain a two-parameter family of geodesics. The union (envelope) of this family defines an emergent two-dimensional surface in the bulk. Motivated by the analogy with string theory, one can then construct an action for this surface—analogous to the Polyakov action—that captures the dynamics of entanglement entropy as the entangling regions change.
@@
where $\partial_{\alpha^{\prime}}=\left(\partial_{\tau^{\prime}},\partial_{\sigma^{\prime}}\right)$ acts only on the base point $X^{\prime}\left(\tau^{\prime},\sigma^{\prime}\right)$. Both $X$ and $X^{\prime}$ are boundary-determined bulk points fixed by the entangling regions in the CFT. Since the two endpoints $X$ and $X^{\prime}$ of any geodesic are interchangeable, there exists a corresponding set of derivatives $\partial_{\alpha}=\left(\partial_{\tau},\partial_{\sigma}\right)$ acting on the field point $X\left(\tau,\sigma\right)$. As a result, we obtain a non-local action constructed from the entanglement entropy, which can describe the physics at either of the spacetime coordinates $X$ or $X^{\prime}$. To verify the consistency of this action, one must check whether it reproduces the entanglement entropy or geodesic length. Before this verification, however, let us examine its physical meaning. Here, instead of treating $X^{\mu}\left(\tau,\sigma\right)$ as the embedding of a true string, the embedding here is constructed out of geodesics in AdS that are dual to entangling intervals in the CFT2. The Synge’s world function $\Omega\left(X,X^{\prime}\right)$ (half the squared geodesic distance) replaces the role of $X^{\mu}$ as the basic dynamical variable. Consequently, the action is built directly from entanglement entropy data (through $\Omega$) rather than fundamental fields. The coordinates $\left(\tau,\sigma\right)$ parametrize the varying entangling regions. As these regions evolve, their dual geodesics sweep out a two-dimensional surface in AdS. This surface is not a fundamental string worldsheet but an effective entanglement worldsheet: a bookkeeping device that tracks the dynamics of entanglement entropy. The action (40) therefore encodes the dynamics of geodesic lengths (equivalently entanglement entropies) as the entangling regions change. The quadratic form in derivatives of $\Omega^{\mu^{\prime}}$ resembles the Polyakov action, ensuring that the dynamics is governed by variations of geodesic distance.
@@
where $g_{\mathrm{ab}}\mathbb{X}^{\mathrm{a}}\mathbb{X}^{\mathrm{b}}$ denotes the squared geodesic distance between the base point $X^{\prime}$ and the field point $X$. A key feature of RNC is that they shift the base point $X^{\prime}$ to the origin, so that $\mathbb{X}^{\mathrm{a}}=0$. On the other hand, since the Synge’s world function $\Omega\left(X,X^{\prime}\right)$ depends on both the base point $X^{\prime}$ and the field point $X$, and since these points can be parameterized by the boundary CFT2 coordinate $\left(\tau,\sigma;\tau^{\prime},\sigma^{\prime}\right)$, we can write $\Omega\left(X,X^{\prime}\right)=\Omega\left(X\left(\tau,\sigma\right),X^{\prime}\left(\tau^{\prime},\sigma^{\prime}\right)\right)$. Consequently, the RNC coordinates $\mathbb{X}^{\mathrm{a}}$ can also be expressed in terms of the CFT2 coordinate $\left(\tau,\sigma;\tau^{\prime},\sigma^{\prime}\right)$ as
@@
$+\frac{1}{\ell^{2}}\int_{\partial\Sigma}d\tau^{\prime}\left[2n_{\alpha^{\prime}}\Omega_{\mu^{\prime}\nu^{\prime}}\partial_{\rho^{\prime}}\Omega^{\mu^{\prime}}\partial^{\alpha^{\prime}}\Omega^{\nu^{\prime}}\delta X^{\rho^{\prime}}\right]_{\sigma=\mathrm{bdy}},$
@@
which is precisely the geodesic equation for the endpoints along the D-brane. This result can be interpreted as shown in figure (7). In our setup, the spacetime point $X$, representing a geodesic endpoint, is parameterized by two-dimensional coordinates $\left(\tau,\sigma\right)$. Consequently, the geodesic sweeps out a two-dimensional surface whose dynamics are governed by the non-local EOM and boundary conditions (50). Taking the near-coincidence limit $X\rightarrow X^{\prime}$ reduces these to the local EOM and boundary conditions (52), in full agreement with string theory. Under the appropriate boundary conditions, the boundary $\partial\Sigma$ obeys the geodesic EOM (53). Thus, the proposed action (40) indeed contains the expected geodesic dynamics.
Figure 7: The left panel shows how a family of geodesics between $X$ and $X^{\prime}$ (red lines) sweeps out an effective two-dimensional surface. The length of each geodesic corresponds to the bulk dual of a CFT entanglement entropy. In the near-coincidence limit $X\rightarrow X^{\prime}$, this surface acquires the structure of a string worldsheet, with the Polyakov action emerging naturally. The right panel illustrates this worldsheet. Depending on the gauge choice, the open string (blue lines) can fluctuate along horizontal or vertical directions, while its endpoints obey the geodesic equation for a point-like particle.
@@
$S_{E}=\frac{1}{\ell^{2}}\int d^{2}\sigma^{\prime}\sqrt{\gamma}\left(\gamma^{\alpha^{\prime}\beta^{\prime}}\Omega_{\mu^{\prime}\nu^{\prime}}\left(X,X^{\prime}\right)+i\epsilon^{\alpha^{\prime}\beta^{\prime}}\mathcal{B}_{\mu^{\prime}\nu^{\prime}}\left(X,X^{\prime}\right)\right)\partial_{\alpha^{\prime}}\Omega^{\mu^{\prime}}\left(X,X^{\prime}\right)\partial_{\beta^{\prime}}\Omega^{\nu^{\prime}}\left(X,X^{\prime}\right).$ (56)
@@
$\mathcal{B}_{\mu^{\prime}\nu^{\prime}}\left(X^{\prime}\right)$ $=$ $B_{\mu^{\prime}\nu^{\prime}}\left(X^{\prime}\right)+\nabla_{\lambda^{\prime}}B_{\mu^{\prime}\nu^{\prime}}\Omega^{\lambda^{\prime}}\left(X,X^{\prime}\right)+\frac{1}{2}\left(\nabla_{\alpha^{\prime}}\nabla_{\beta^{\prime}}B_{\mu^{\prime}\nu^{\prime}}\left(X^{\prime}\right)-\frac{1}{3}R_{\;\alpha^{\prime}\mu^{\prime}\beta^{\prime}}^{\lambda^{\prime}}B_{\lambda^{\prime}\nu^{\prime}}\right$. (57)